The Global Growth Fund is a load fund with a 3 percent front load fee. It started the year with a net asset value (NAV) of $17.00. During the year the fund distributed $1.30, and at the end of the year its NAV was $18.45. What was the fund's return, and what was an investor's return? Round your answers to two decimal places.

The fund's return:   %

Return on the individual's investment:   %

Why are they different?

The answers are different because the fund reports its return based on the -Select-purchase pricenet asset valueItem 3 and the investrors calculate their return based on the -Select-purchase pricenet asset valueItem 4 .
Answer #1

Details Provided : Front load fee = 3%, Start NAV = 17, Funds Distributed = 1.3, End of Year NAV = 18.45

Suppose an investor puts in 100 in this funds

NAV of customer = (Invested amount - load fee) / Start NAV = (100-3)/17 = 5.7059

Amount at the end of year = 5.7059 * 18.45 = 105.27

Funds return = (End NAV - Start NAV + Dividend Distributed ) / Start NAV = 16.18%

Individual investment Return = 105.27 - 100 + 1.3 = 8.27%

They are different as there is a front load of fees charged by fund house